What is the maturity value if the principle is $1400.00, and the rate is 16% compounded quarterly over 6 yrs?
Please help,
Ellen
: What is the maturity value if the principle is $1400.00, and the rate is 16% compounded quarterly over 6 yrs?
Please help,
Ellen

Answer by scott8148(2719)

16% quarterly is 4% (16/4) per compounding __ 6 years is 24 quarters (compoundings)

A=P(1+r)^n __ MV=1400(1+.04)^24 __ MV=3588.63 (approx)